OpenStudy (anonymous):

Henry Devine bought a new dishwasher for $320. He paid $20 down and made 10 monthly payments of $34. What actual yearly interest rate did Henry pay? A. 29.09 B. 14.55 C. 68.75 D. 34.38

OpenStudy (anonymous):

I did the math 2 times and I got the answer B twice, but when I submitted my test. Answer B was wrong, I've looked all over the internet for help and I can't seem to find the answer. Someone please help me.

OpenStudy (redohawk):

i'm thinking D

OpenStudy (anonymous):

Can you try to explain to me how you came up with that answer please @RedoHawk

OpenStudy (anonymous):

Down Payment= $20 10 Payments=10x34, =340 Total paid=340+20=$360 Cost price was........$320 Interest=.................._____40 The formula for Simple Interest is i=Prt, where i=annual interest rate P=starting amount r= interest rate, as a percent t=time, in years From i=Prt, r=i/Pt i=40, P=320, t=10/12 of a year r=40/320(10/12) r=40/266.67 r=0.1499981 r=15% @Tayy-Renae1317